# On almost commuting unitary matrices

### v1

@@ Abstract.
A question going back to Halmos asks when two approximately commuting matrices of a certain kind are close to exactly commuting matrices of the same kind. It has long been known that there is a winding number obstruction for approximately commuting unitary matrices to be close, in a dimension-independent way, to genuinely commuting unitary matrices. In this paper, under the vanishing of the said obstruction, we obtain effective bounds for the distance to commuting unitary matrices in terms of the commutator of the original matrices.
A. D. was partially supported by NSF / BSF grants 2350543 / 2023695 (respectively), 2452324 / 2024734 (respectively) and a DFG Middle-Eastern collaboration project 529300231.
L. H. was supported as a Zuckerman Postdoctoral Scholar in the 2023-2024 cohort.
@@ 1. Introduction
By considering $a=x+iy$, one can restate the self-adjoint version of the above question as whether the distance from $a$ to normal operators can be estimated in terms of $\|[a,a^{*}]\|$ (that is, the norm of the self-commutator of $a$). In this form, positive answer to the question was provided in [22]. See also [13] for a shorter proof. In both cases, it is reduced by contradiction to a lifting problem in a certain ultraproduct of matrix algebras, and extracting an explicit dependence relation between $\varepsilon$ and $\delta$ from this kind of proof becomes a challenging reverse engineering problem with unclear possibilities of obtaining a optimal bound. An alternative approach to the self-adjoint problem was proposed in [21], which led to an order sharp estimate $\operatorname{dist}(a,\textrm{\{normal matrices\}})\leq C\|[a,a^{*}]\|^{1/2}$, corresponding to $\delta=C\varepsilon^{2}$. See also [17, 19] regarding earlier results and [8] for a survey.
@@ 1.1. Almost commuting unitary matrices: overview and the main result
We will be mainly interested in Question 1.1 for a pair of unitary matrices (which will be denoted by $u$ and $v$ rather than $x$ and $y$). In this case, the problem presents a topological obstruction, see [25, 12, 11], which has been characterized in several somewhat equivalent ways, the most simple of them being the winding number invariant: for $u,v\in U(n)$ with $\|[u,v]\|<2$, the winding number $w(u,v)$ of the curve
@@
is well-defined. Assuming that $\varepsilon$ is small enough, it is shown in the above references that any pair $u,v\in U(n)$ that is close to a commuting pair must have $w(u,v)=0$. This is confirmed by the example of Voiculescu’s unitaries [25], a family of pairs of matrices $u_{n},v_{n}\in U(n)$ with $\|[u_{n},v_{n}]\|=O(1/n)$ whose distance to any commuting pair is bounded from below by $\sqrt{2}-1-o(1)$ as $n\to+\infty$ [11].
In view of the discussion above, a natural setting for Question 1.1 for two unitaries includes the additional assumption that their winding number invariant vanishes. In this form, [10] and [15] independently resolved the problem around the same time. Similarly to the self-adjoint case considered in [22, 13], both solutions involve the ultraproduct construction and therefore do not provide an explicit relation between $\varepsilon$} and $\delta$.
The present paper answers Question 1.1 for unitary matrices in the above setting with the following quantitative bounds:
@@
There exists an absolute constant $C>0$ such that for any $u,v\in U(n)$ with $w(u,v)=0$ one can find $u^{\prime},v^{\prime}\in U(n)$ such that
$u^{\prime},v^{\prime}]=0,\quad\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq C\|[u,v]\|^{1/30}.$
1.2. Overview of the proof
First, we note that there are three versions of the winding number invariant, with different potential directions of generalizations; however, all three coincide in our setting of a pair of almost commuting unitary matrices. We will use two of them: the winding number $w(u,v)$ of the curve defined by equation (1.3) and the isospectral invariant $\operatorname{isospec}(u,v)$, originally introduced in [5, Theorem 4.1], see equation (3.2) in Section 3. For the convenience of the reader, we show that $w(u,v)=\operatorname{isospec}(u,v)$ in Corollary 3.4, since the original proof is spread among several of the above-referenced papers.
The proof of Theorem 1.2 is a combination of the following two results obtained in this paper. The first one provides quantitative estimates for the isospectral homotopy lemma from [5].
@@ 1.2. Overview of the proof
There exists an absolute constant $C>0$ such that for every two unitary matrices $u,v\in U(n)$ with $\operatorname{isospec}(u,v)=0$, there exists a piecewise smooth path $v_{t}\colon[0,1]\to U(n)$ of fixed length with
@@
The second result gives a quantitative variant of a theorem, whose analogue in the setting of real rank zero $C^{*}$-algebras was proved in [23] and [24] independently and around the same time.
@@
Suppose that $u,v$ are unitary matrices such that $\|[u,v]\|\leq\delta$ and there exists a continuous path $\{u_{t}\colon t\in[0,1]\}$ such that:
@@
Proof of Theorem 1.2. As mentioned earlier, Corollary 3.4 implies that $\operatorname{isospec}(u,v)=0$. Then, the result follows from Theorem 1.4 applied with $\delta=C\|[u,v]\|^{2/5}$, using the conclusion of Lemma 1.3.  ∎
In the cases of Lemma 1.3 and Theorem 1.4, we focus on the setting of two unitary matrices. It is likely that Lemma 1.3 and 1.4 can be proved in the level of generality close to the original results [5, 23]. We intend to explore it in the future, as well as potential generalizations of Theorem 1.2 beyond the matrix case.
The proof of Lemma 1.3, is obtained by retracing the steps in [5]. While the direct repetition of the arguments would not necessarily lead to the exponent $2/5$, one can make certain optimizations along the way, see Subsection 3.2.
Theorem 1.2 is more difficult. While it does not technically involve $C^{*}$-algebras, we believe that this viewpoint best explains our motivation. Note that, if one of the unitaries has a large spectral gap, then the problem becomes “topologically trivial” and, essentially, reduces to the main result of [21], see Proposition 4.1 and Lemma 4.2. Our construction of commuting approximants, which involves several applications of the above “gapped” result, involves a sequence of operator-theoretic procedures (mainly applied to the matrix $u$), and at each step one needs to maintain control over the norm of the commutator with $v$. Not every operator-theoretic procedure respects this property: for example, considering spectral projections or discontinuous functional calculus may blow up the norm of the commutator.
Difficulties of similar kind appeared in the proof of the Brown–Pedersen Conjecture in [23] (see also [24]). In this case, the process could also be described as gap opening by a sequence of operator-theoretic procedures, which have to be performed inside of a particular $C^{*}$-algebra. Again, not all operator-theoretic procedures are $C^{*}$-algebraic”, with the most simple example being the consideration of arbitrary spectral projections.
Our main observation, which appeared in [21] in a somewhat different context, is that many $C^{*}$-algebraic” procedures also have “commutator control” properties. For example, smooth functional calculus for normal elements has this property due to theory of operator Lipschitz functions, see (OL4) and (OL6) in the Appendix. The proof in Section 4.2 is the result of such re-engineering of the whole construction in [23]. While things such as smooth calculus can be considered straightforward, we would like to draw the reader’s attention to the following two much less trivial aspects.
@@
    The main result of [23] states that one can open a gap in a unitary element of a real rank zero $C^{*}$-algebra $\mathcal{A}$, provided that it belongs to the connected component of the identity in the unitary group. The isospectral homotopy lemma 1.3 is a quantitative commutator analogue of this assumption.
@@
    The real rank zero property itself states that every self-adjoint element of the $C^{*}$-algebra can be approximated by elements with finite spectra. One can restate it as follows: every self-adjoint element $a\in\mathcal{A}$ is close to another self-adjoint element $a^{\prime}$, such that taking spectral projections of $a^{\prime}$ is a valid $C^{*}$-algebraic operation. The commutator analogue of this property is Claim (4) of Proposition 4.1 (which relies the main results of [21]). Indeed, if $a=a^{*}$ and $[a,v]$ is small, one can apply the claim once and produce an exactly commuting pair $a^{\prime},v^{\prime}]=0$. Afterwards, if $f$ is a bounded Borel function, then $f(a^{\prime}),v^{\prime}]=0$. Since $v$ is close to $v^{\prime}$, the commutator $f(a^{\prime}),v]$ will remain small.
We note that each step of this kind leads to some losses in the commutator norm, and different interpretations of abstract $C^{*}$-algebraic constructions can lead to vastly different quantitative results. Controlling these losses was one of the main technical challenges the proof.
@@ 1.3. Structure of the paper
In Section 2, we establish some (mostly well-known) preliminaries from spectral theory, involving estimates of norms of spectral projections associated to disjoint intervals in Subsection 2.1, some general properties of projections and unitary operators in Subsection 2.2, and a version of a more specific result on intertwined families of projections, obtained in [5] in its original form, in Subsection 2.3. In Section 3, we introduce the isospectral invariant in Subsection 3.1, prove Lemma 1.3 (the isospectral homotopy lemma) in Subsection 3.2, and use it to establish equivalence between two invariants in Subsection 3.3. In Section 4, which is the most technical in the paper, we prove Theorem 1.4. Subsection 4.1 contains some preparations, including Lemma 4.2 on almost commuting unitaries with a spectral gap. Following the general ideas in [23], we use Proposition 4.3 to construct an amplification of the original almost commuting pair and perturb it in the larger space to create a spectral gap in Subsection 4.2. This allows to find commuting approximants in a larger space in Subsection 4.3. Afterwards, one needs to descend back into the original space, which is done in two steps, in Subsections 4.4 and 4.5.
@@ 2. Preliminaries
In this section, we gather several (mostly already known) results regarding spectral projections, approximate spectral projections, and behavior of spectral projections under small perturbations. The proofs are included mostly for the sake of completeness, as well as in order to introduce notation that will be useful in the later sections.
We will actively use the results from the theory of operator Lipschitz functions, often citing properties (OL1) – (OL8) stated in the Appendix. Let $u_{1},u_{2}\in U(n)$ and $g\colon\mathbb{T}=\mathbb{R}/\mathbb{Z}\to\mathbb{R}$ be a smooth function on the circle. Then $\|g(u_{1})-g(u_{2})\|$ is small provided that $\|u_{1}-u_{2}\|$ is small. As a consequence, if $f$ is another bounded function whose support is disjoint from that of $g$, we have $f(u_{1})g(u_{1})=0$ and therefore $\|f(u_{1})g(u_{2})\|$ is small.
If $g$ is, say, discontinuous, then the first claim from the previous paragraph falls apart. However, some estimates of this kind remain valid if the supports of $f$ and $g$ are sufficiently disjoint. Most of the results below have been used in previous work such as [5, 9], in many cases with better constants. The use of operator Lipschitz functions allows us to provide the proofs in a somewhat unified language. We begin with an elementary fact about bump-type functions.
@@
From (OL7) in the Appendix, the problem can be restated in the language of $1$-periodic function on $\mathbb{R}$. By rescaling (OL5), at the expense of the factor $1/\beta$ one can consider $\beta=1$. In this case, one can construct $f$ as a linear combination of bump functions with disjoint supports which implies a uniform bound in $C_{b}^{2}(\mathbb{R})$ and allows to complete the proof by applying (OL3). ∎
@@ 2.1. Behavior of spectral projections under small perturbations
The following two lemmas, used in [5, Lemma 2.2 & Lemma 2.9] and other related work, formalize the principle that the orthogonality relations between spectral projections (and, more generally, functions with disjoint supports) are approximately preserved under small perturbations, provided that the boundaries of the intervals under consideration are separated from one another.
@@
Let $g_{j}\colon\mathbb{T}\to[-1,1]$ be measurable and $u_{j}\in U(n)$, $j=1,2$.
@@
$h_{1}|_{I}\equiv 1,\quad h_{2}|_{J}\equiv 0;\quad h_{1}|_{J}\equiv 0,\quad h_{2}|_{I}\equiv 1,\quad\mathrm{supp}\,h_{1}\cap\mathrm{supp}\,h_{2}=\varnothing.$
@@
which implies part (i). Here, we used the fact that $h_{1}(u)h_{2}(u)=(h_{1}h_{2})(u)=0$ and the definition of $\|\cdot\|_{\operatorname{OL}}$ (see Appendix). Part (ii) follows from Part (i) applied to the functions $1-g_{2}$ and $g_{1}$. ∎
@@
Let $I,J\subset\mathbb{T}$ be two intervals with disjoint boundaries, and $u_{j}\in U(n)$, $j=1,2$. Then
@@
2.2. Some general results on orthogonal projections and unitary operators

### v2

@@ Abstract.
A question going back to Halmos asks when two approximately commuting matrices of a certain kind are close to exactly commuting matrices of the same kind. It has long been known that there is an obstruction for approximately commuting unitary matrices to be close, in a dimension-independent way, to genuinely commuting unitary matrices. In this paper, under the vanishing of the said obstruction, we obtain effective bounds for the distance to commuting unitary matrices in terms of the commutator of the original matrices.
A. D. was partially supported by NSF / BSF grants 2350543 / 2023695 (respectively), 2452324 / 2024734 (respectively), a Horizon Marie-Curie SE project no. 101086394, and a DFG Middle-Eastern collaboration project 529300231.
L. H. was supported by a Zuckerman Postdoctoral Fellowship in the 2023-2024 cohort.
@@ 1. Introduction
By considering $a=x+iy$, one can restate the self-adjoint version of the above question as whether the distance from $a$ to the set $\mathcal{N}$ of normal matrices can be estimated in terms of $\|[a,a^{*}]\|$ (that is, the norm of the self-commutator of $a$). In this form, a positive answer to the question was provided in [22]; see also [13] for a shorter proof. In both cases, the proof starts by assuming towards contradiction, and is reduced to a lifting problem in a certain asymptotic matrix sequence $C^{*}$-algebra. Extracting an explicit dependence relation between $\varepsilon$ and $\delta$ from this kind of proof then becomes a challenging reverse engineering problem, and it is unclear whether the bounds obtained in this way are a optimal. An alternative approach to the self-adjoint problem was proposed in [21], and has led to an order sharp estimate $\operatorname{dist}(a,\mathcal{N})\leq C\|[a,a^{*}]\|^{1/2}$, corresponding to $\delta=C\varepsilon^{2}$. See also [17, 19] regarding earlier results and [8] for a survey.
@@ 1.1. Almost commuting unitary matrices: overview and the main result
We will be mainly interested in Question 1.1 for a pair of unitary matrices. In this case, the problem presents a topological obstruction, see [28, 12, 11], as well as the paragraph following [8, Question I.5]. This obstruction has been characterized in several somewhat equivalent ways, one of the most simple of them being the so-called winding number invariant. For $u,v\in{\mathrm{U}}(n)$ with $\|[u,v]\|<2$, we denote by $w(u,v)$ the winding number (with respect to the origin) of the curve
@@
Assuming that $\varepsilon$ is small enough, it is shown in the above references that any pair $u,v\in{\mathrm{U}}(n)$ that is $\varepsilon$-close to a commuting pair must have $w(u,v)=0$. The winding number is indeed an obstruction due to the example of Voiculescu’s unitaries [28], which is a sequence of pairs of unitary matrices $u_{n},v_{n}\in{\mathrm{U}}(n)$ satisfying $\|[u_{n},v_{n}]\|=O(1/n)$ while $w(u_{n},v_{n})=-1$ [11], thereby showing that $u_{n},v_{n}$ are not eventually close to a commuting pair of matrices.
In view of the discussion above, a natural setting for Question 1.1 for two unitaries includes the additional assumption that their winding number vanishes. In this stated form, [10] and [15] independently resolved the problem at around the same time. Similarly to the self-adjoint case considered in [22, 13], both solutions involve an asymptotic sequence algebra construction, and therefore do not provide an explicit relation between $\varepsilon$} and $\delta$, leaving the corresponding quantitative version of Question 1.1 unresolved, as pointed out in the discussion preceding [18, Theorem 4.10].
The main result of the present paper answers the above question by providing quantitative bounds for the distance in terms of the norm of the commutator for unitary matrices with vanishing winding number obstruction as follows:
@@
There exists an absolute constant $C>0$ such that for any $u,v\in{\mathrm{U}}(n)$ with $w(u,v)=0$ one can find a pair $u^{\prime},v^{\prime}\in{\mathrm{U}}(n)$ such that
$\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq C\|[u,v]\|^{1/30},\quad[u^{\prime},v^{\prime}]=0.$
We next provide an overview of the proof. We start by noting that there are various versions of the obstruction to proximity to a nearly commuting pair, with different potential directions of generalizations; however, they all coincide in our setting of pairs of almost commuting unitary matrices. We will use the winding number $w(u,v)$ of the curve defined by equation (1.3) and the isospectral invariant $\operatorname{isospec}(u,v)$, originally introduced in [5, Theorem 4.1], see equation (3.2) in Section 3. For the convenience of the reader, we show that $w(u,v)=\operatorname{isospec}(u,v)$ in Corollary 3.4, since the original proof is spread among several of the above-referenced papers. The proof of Theorem 1.2 is a combination of the following two results. The first one is a quantitative version of the isospectral homotopy lemma from [5].
@@
There exists an absolute constant $C>0$ such that for every two unitary matrices $u,v\in{\mathrm{U}}(n)$ with $\operatorname{isospec}(u,v)=0$, there exists a piecewise smooth path $v_{t}\colon[0,1]\to{\mathrm{U}}(n)$ of fixed length with
@@
The second result is the following quantitative variant of a theorem, whose analogue in the setting of real rank zero $C^{*}$-algebras was proved in [24] and [26] independently and at around the same time.
@@
Suppose that $u,v$ are unitary matrices such that $\|[u,v]\|\leq\delta$ and that there exists a continuous path $\{u_{t}\colon t\in[0,1]\}$ such that:
@@
Proof of Theorem 1.2. Since $w(u,v)=0$, Corollary 3.4 implies $\operatorname{isospec}(u,v)=0$. Then, the result follows from Theorem 1.4 applied with $\delta=C\|[u,v]\|^{2/5}$, using the conclusion of Lemma 1.3.  ∎
The proof of Lemma 1.3 is obtained by retracing and providing quantitative bounds in several steps in [5]. While the direct repetition of the arguments does not lead to the exponent $2/5$, we make certain optimizations along the way to acheive this (see Subsection 3.2).
Theorem 1.4 presents a different kind of challenge, and, while it does not technically involve $C^{*}$-algebras, the corresponding language of $C^{*}$-algebras best explains our motivation. We first note that, if one of the unitaries has a large gap in its spectrum, then the problem becomes “topologically trivial” and essentially reduces to the main result of [21] (see Proposition 4.1, Lemma 4.3, and Remark 4.7). Our construction of commuting approximants, which involves several applications of the above “gap-opening” result, involves a sequence of operator-theoretic procedures, mainly applied to the matrix $u$, where at each step one needs to maintain control over the norm of the commutator with $v$. Not every operator-theoretic procedure has this property. For example, considering spectral projections or discontinuous functional calculus may blow up the norm of the commutator.
Difficulties of a similar kind appeared in the proof of the (first) Brown–Pedersen Conjecture in [24] and [26], which asks whether the multiplier algebra of a real-rank zero $C^{*}$-algebra remains of real-rank zero. In that context, the process could also be described as gap-opening by a sequence of operator-theoretic procedures, to be performed while remaining inside of the given $C^{*}$-algebra.
Our main observation, which was also used in [21], is that many procedures that allow us to remain in a given $C^{*}$-algebra also have “commutator control” properties. For example, smooth functional calculus for normal elements has this property due to theory of operator Lipschitz functions (see (OL4) and (OL6) in Appendix A). Therefore, the aspects of certain $C^{*}$-algebraic proofs can be transferred into proofs that involve almost commuting matrices with commutator control. The proof in Subsection 4.2 is the result of such a transfer of the construction in [24]. While smooth functional calculus may be considered straightforward, we draw the reader’s attention to the following three more involved aspects of our work.
@@
    The main result of [24] states that one can open a gap in a unitary element of a real rank zero $C^{*}$-algebra, provided that it belongs to the connected component of the identity in the unitary group. In Lemma 1.3 we provide a quantitative refinement of the isospectral homotopy lemma from [5, Lemma 6.1], which connects one unitary to the identity with controlled losses to commutator control along the path. The transfer from path connectedness to the identity into commutator control properties suggests a quantitative refinement of [24], which we realize in Theorem 1.4.
@@
    The real-rank zero property for a given $C^{*}$-algebra $\mathcal{A}$ is equivalently stated as requiring that every self-adjoint element $a$ in $\mathcal{A}$ is close to another self-adjoint element $a^{\prime}$ in $\mathcal{A}$ whose spectral projections remain in the $C^{*}$-algebra $\mathcal{A}$. Item (4) of Proposition 4.1 captures the commutator analogue of this property. Indeed, if $a=a^{*}$ and $[a,v]$ is small, one can apply item (4) of Proposition 4.1 once, and produce an exactly commuting pair $a^{\prime},v^{\prime}]=0$. Afterwards, if $f$ is a bounded Borel function, then $f(a^{\prime}),v^{\prime}]=0$. Since $v$ is close to $v^{\prime}$, the commutator $f(a^{\prime}),v]$ will therefore remain small.
  3. (3)
    Most of the constructions from [24] happen in the amplification of the original $C^{*}$-algebra, and it is an important fact that any amplification of a real rank zero algebra is still of real rank zero. The commutator analogue of this, considered in the previous paragraph, is also preserved upon amplification, but with an increasingly large losses in the estimates as the size of the amplifications increases. In that approach, the estimates would also depend on the length of the path obtained in Lemma 1.3. Instead of this, the strategy in our paper is to use [21] in the amplified matrix algebra directly, thereby avoiding an (exponential) loss in the final distance bounds.
We note that each “transfer” of an abstract $C^{*}$-algebraic construction into the commutator-controlled language leads to some loss in the exponent of the commutator norm in the main result, and different interpretations of these constructions can lead to vastly different quantitative results. Controlling these losses was one of the main technical challenges in our work.
@@ 1.2. Structure of the paper
In Section 2, we state some (mostly known) results on spectral projections and their behavior under small perturbations. In Subsection 2.1, we discuss various relations between spectral projections associated to pairs of intervals with disjoint endpoints. In Subsection 2.2, we include some general properties of projections and unitary operators. Subsection 2.3 deals with a quantitative version of the result on intertwined families of projections, whose qualitative analogues were obtained in [5].
In Section 3, we discuss our topological invariants for pairs of unitary matrices. In Subsection 3.1, we discuss a simplified version of the isospectral invariant. In Subsection 3.2, we prove Lemma 1.3m which is the isospectral homotopy lemma with commutator control. In 3.3, our isospectral homotopy lemma is used to establish an equivalence between the winding number and isospectral invariants. In Subsection 3.4, we discuss an equivalent definition of the winding number invariant from [12], as well as situations where there is no obstruction to proximity to a commuting unitary pair in infinite dimensions. In Section 4 we prove Theorem 1.4. Subsection 4.1 contains some preparations, including Lemma 4.3 on almost commuting unitary matrices with one having a spectral gap. Following and transferring the general ideas of [24] to the commutator-controlled setting, we use Proposition 4.5 to construct an amplification of the original almost commuting pair and perturb it in the larger space to create a spectral gap in Subsection 4.2. This allows to find commuting approximants in a larger space in Subsection 4.3. Afterwards, we descend these approximants back into the original space, which is done in two dimension-reduction steps in Subsections 4.4 and 4.5.
@@ 2. Preliminaries
In this section, we gather several (mostly well known) results on spectral projections associated to isolated intervals in the spectra of unitary matrices, and their behavior under small perturbations. The proofs are included mostly for the sake of completeness, as well as to introduce notation that will be useful in the later sections.
We will actively use the results from the theory of operator Lipschitz functions, often citing properties (OL1) – (OL8) stated in Appendix A. Let $u_{1},u_{2}\in{\mathrm{U}}(n)$ and $g\colon\mathbb{T}=\mathbb{R}/\mathbb{Z}\to\mathbb{R}$ be a smooth function on the circle. Then $\|g(u_{1})-g(u_{2})\|$ is small provided that $\|u_{1}-u_{2}\|$ is small. As a consequence, if $f$ is another bounded function whose support is disjoint from that of $g$, we have $f(u_{1})g(u_{1})=0$ and therefore $\|f(u_{1})g(u_{2})\|$ is small.
If $g$ is, say, discontinuous, then the first claim from the previous paragraph falls apart. However, some statements of this kind remain true if the supports of $f$ and $g$ are sufficiently disjoint. Most of the results below have been used in previous work such as [5, 9], in many cases with better constants. The use of operator Lipschitz functions allows us to provide the proofs in a somewhat unified language. We begin with an elementary fact about bump-type functions.
@@
From (OL7) in Appendix A, the problem can be restated in the language of $1$-periodic function on $\mathbb{R}$. By rescaling (OL5), at the expense of the factor $1/\beta$ one can consider $\beta=1$. In this case, one can construct $f$ as a linear combination of bump functions with disjoint supports which implies a uniform bound in $C_{b}^{2}(\mathbb{R})$ and allows to complete the proof by applying (OL3). ∎
@@ 2.1. Behavior of spectral projections under small perturbations
The following two lemmas, used in [5, Lemma 2.2 & Lemma 2.9] and other related work, formalize the principle that the orthogonality relations between spectral projections (and, more generally, functions with disjoint supports) are approximately preserved under small perturbations, provided that the boundaries of the intervals under consideration are sufficiently separated from one another.
@@
Let $g_{j}\colon\mathbb{T}\to[-1,1]$ be measurable and $u_{j}\in{\mathrm{U}}(n)$, $j=1,2$.
@@
$h_{1}|_{I}\equiv 1,\quad h_{2}|_{J}\equiv 0;\quad h_{1}|_{J}\equiv 0,\quad h_{2}|_{I}\equiv 1,\quad\mathrm{supp}\,h_{1}\cap\mathrm{supp}\,h_{2}=\emptyset.$
@@
which implies part (i). Here, we used the fact that $h_{1}(u)h_{2}(u)=(h_{1}h_{2})(u)=0$ and the definition of $\|\cdot\|_{\operatorname{OL}}$ (see Appendix A). Part (ii) follows from Part (i) applied to the functions $1-g_{2}$ and $g_{1}$. ∎
@@
Let $I,J\subset\mathbb{T}$ be two intervals with disjoint boundaries, and $u_{j}\in{\mathrm{U}}(n)$, $j=1,2$. Then
@@
2.2. Perturbations of orthogonal projections and unitary operators